# Analyzing and Mitigating Bias for Vulnerable Classes: Towards Balanced Representation in Dataset

### v1

@@ III-E Bias Assessment and Model Re-calibration
Model Re-calibration: This involves dynamically adjusting class weights within the loss function, guided by real-time performance metrics. This ensures the model stays optimized to any shifts in class representation or emerging imbalances. When new class samples or data become available, the model undergoes re-training to stay aligned with the evolving operational context. The last step further includes an iterative refinement, monitoring and updating bias mitigation strategies based on the latest performance assessments and re-calibrations.
@@ IV Experimental Evaluation
IV-A Test Setup
Test Set: The overall dataset comprises approximately 1.4 million images, with annotations across several classes. Our focus is on the Pedestrian (149,921), Cyclist(17,060), and Motorcyclist (16,779) classes, as it provides a rich source for analyzing biases in object detection.
@@ IV-A Test Setup
IV-B Hyperparameter Optimization
Initial Hyperparameters: Were set as follows: learning rate of 0.001, batch size of 32, and weight decay of 0.0001 for CNN models. For ViT we configured the training with a batch size of 32 and an initial learning rate of 1e-3, which follows a linear decay to 1e-5. The model was trained for 30 epochs using the Adam optimizer. A dropout rate of 0.1 and a weight decay of 0.03 were applied to avoid overfitting.
@@ IV-B Hyperparameter Optimization
IV-C Training Process
Training Details: As our sample size is small, CNN models were trained for 50 epochs using an Adam optimizer. A learning rate scheduler was used to reduce the learning rate by 10% every 10 epochs. Regularization techniques such as dropout and data augmentation (rotations and flipping) were used to prevent overfitting [16].
Data Splits: The dataset is divided into 70% training, 15% validation, and 15% test splits, ensuring a diverse representation of scenarios in each set.
Computational Resources: Training is conducted on a high-performance computing cluster with a combination of NVIDIA Tesla V100 GPUs using Open MPI and PyTorch as the deep learning framework.
IV-D Model Evaluation
Performance Metrics: Models were evaluated based on mAP, %IoU and NDS (Intersection over Union, NuScenes Detection Score), alongside sensitivity and selectivity scores for class-specific analysis.
@@ IV-D Model Evaluation
Class-specific Analysis: Performance metrics for each class were analyzed to assess improvements in detecting pedestrians, cyclists, and motorcyclists. Attention was paid to false positive or negative rate changes for these classes.
IV-E Bias Impact Analysis
Error Analysis: Figure 2 shows error analysis for the discussed classes. The ResNet18 model exhibited higher false positive rates across all classes than the ViT model, indicating a tendency towards over-prediction. Conversely, the ViT model maintained lower false positives but demonstrated higher false negatives, particularly for Pedestrians and Cyclists, suggesting a cautious but potentially under-detecting approach. Notably, the ViT model was more balanced for the Motorcyclist class, with significantly lower false negatives than ResNet18, indicating more reliable detection.
@@ IV-E Bias Impact Analysis
Figure 2: Figure shows class-wise error rate for ResNet and ViT according to False Negatives and False Positives in model validation
IV-F Discussion of Results
Figure 3: Figure showing heatmap for all three classes
@@ IV-F Discussion of Results
For measuring IoU, mAP and NDS scores mentioned in Tables 1 and 2, we used a subset approach, where the representation of one class (e.g., pedestrian) is kept at 67% and the other two classes have equal representation. This ratio is adjusted during iterations until all classes have the same representation. The purpose is to capture metrics in a scenario where the weighted class functions are normalized. Table I shows the Intersection over Union (%IoU) for different classes and the NuScenes Detection Score (NDS) across various conditions for ResNet18. Pedestrian detection remains relatively high in all conditions, with %IoU above 70% except when combined adverse conditions are present. Cyclist detection suffers significantly in all but normal conditions, with %IoU dropping to as low as 7.5% during adverse weather. Motorcyclist detection is also affected by these conditions but to a lesser extent. The overall NDS reflects these trends, with the highest score of 72.0 in normal conditions and the lowest of 31.2 during adverse weather. The rotated condition only slightly impacts the %IoU for pedestrians and motorcyclists but more so for cyclists. Mixed conditions present a challenging scenario with a noticeable decrease in %IoU across all classes and a resultant NDS of 35.3.
@@
Table II shows the mean average precision (mAP) of pedestrian, cyclist, and motorcyclist detection under various conditions. Under normal conditions, pedestrian detection is high at 86.3%, while cyclist detection lags at 54.7%. Motorcyclist detection is reasonably high at 77.8%, leading to an overall mAP of 72.93%. However, performance drops significantly at night and during adverse weather, with the lowest cyclist mAP at 7.5% and 8.3%, respectively. Interestingly, rotating images in normal conditions slightly reduces pedestrian and motorcyclist detection by about 1%, but notably impacts cyclist detection, dropping to 52.3%. The combination of all adverse conditions drastically reduces overall performance to a mAP of 36.7%, showcasing the challenges faced by the models in less-than-ideal scenarios.
@@
IV-G Key Takeaways
@@
  • Post-mitigation performance, which involves resampling approaches, shows improvements in both CNN (ResNet18) and ViT models.

### v2

@@ III-E Bias Assessment and Model Re-calibration
Model Re-calibration: This involves dynamically adjusting class weights within the loss function, guided by real-time performance metrics. The step ensures the model stays optimized for any shifts in class representation or emerging imbalances. When new class samples or data become available, the model undergoes re-training to stay aligned with the evolving operational context. The last step includes an iterative refinement, monitoring and updating bias mitigation strategies based on the latest performance assessments.
@@ IV Experimental Evaluation
This section covers training, testing and evaluation methods. The overall dataset comprises approximately 1.4 million images, with annotations across several classes. Our focus is on the Pedestrian (149,921), Cyclist(17,060), and Motorcyclist (16,779) classes, as it provides a rich source for analyzing biases in object detection.
@@
IV-A Hyperparameter Optimization
Initial hyperparameters were set as follows: learning rate of 0.001, batch size of 32, and weight decay of 0.0001 for CNN models. For ViT we configured the training with a batch size of 32 and an initial learning rate of 1e-3, following a linear decay to 1e-5. The model was trained for 30 epochs using the Adam optimizer. A dropout rate of 0.1 and a weight decay of 0.03 is applied to avoid overfitting.
@@ IV-A Hyperparameter Optimization
IV-B Training Process
Our sample size is small, so CNN models were trained for 50 epochs using an Adam optimizer. A learning rate scheduler was used to reduce the learning rate by 10% every 10 epochs. Regularization techniques such as dropout and data augmentation (rotations and flipping) were used to prevent overfitting [18].
The dataset is divided into 70% training, 15% validation, and 15% test splits, ensuring a fair representation of scenarios in each set. Training is conducted on a high-performance computing cluster with a combination of NVIDIA Tesla V100 GPUs using Open MPI and PyTorch as the deep learning framework.
Figure 3: Figure shows layerwise selectivity for the Classes
IV-C Model Evaluation
Models were evaluated using mAP, %IoU and NDS (Intersection over Union, NuScenes Detection Score), alongside sensitivity and selectivity scores for class analysis.
@@ IV-C Model Evaluation
Class-specific Analysis: Performance metrics for each class were used to assess improvements in detecting pedestrians, cyclists, and motorcyclists. For analysis, attention is given to false positive or negative rate changes for these classes. The tests for CNN can be performed using layer-wise analysis and by varying data diversity where one class has a dominant presence.
Figure 4: Visualization of class-wise error rate for ResNet and ViT according to False Negatives and False Positives in model validation.
Layer-wise Analysis: This analysis reveals whether specific layers are biased towards particular classes, which provides strategies for bias mitigation. For instance, if early layers are biased towards detecting pedestrians more than cyclists or motorcyclists, there will be a need to adjust the training data with inclusiveness.
Analysis over Epochs: Monitoring sensitivity and selectivity over epochs allows you to see if the model improves its ability to recognize underrepresented classes. If sensitivity and selectivity for these classes do not improve, or if they plateau early, this could indicate that the model is not effectively learning features for these classes, suggesting that further intervention is required. A layerwise selectivity for the three classes in different scenarios is shown in Figure 4.
Figure 5: Figure showing heatmap for all three classes.
IV-D Bias Impact Analysis
Error Analysis: Figure 4 shows error analysis for the discussed classes. The ResNet18 model exhibited higher false positive rates across all classes than the ViT model, indicating a tendency towards over-prediction. Conversely, the ViT model maintained lower false positives but demonstrated higher false negatives, particularly for Pedestrians and Cyclists, suggesting a cautious but potentially under-detecting approach. Notably, the ViT model was more balanced for the Motorcyclist class, with much lower false negatives than ResNet18, indicating more reliable detection.
@@ IV-D Bias Impact Analysis
IV-E Discussion of Results
For measuring IoU, mAP and NDS scores mentioned in Tables 1 and 2, we used a subset approach, where the representation of one class (e.g., pedestrian) is kept at 67% and the other two classes have equal representation. This ratio is adjusted during iterations until all classes have the same representation. The purpose is to capture metrics in a scenario where the weighted class functions are normalized. Table I shows the Intersection over Union (%IoU) for different classes and the NuScenes Detection Score (NDS) across various conditions for ResNet18. Pedestrian detection remains relatively high in all conditions, with %IoU above 70% except when combined adverse conditions are present. Cyclist detection suffers significantly in all but normal conditions, with %IoU dropping to as low as 7.5% during adverse weather. Motorcyclist detection is also affected by these conditions but to a lesser extent. The overall NDS reflects these trends, with the highest score of 72.0 in normal conditions and the lowest of 31.2 during adverse weather. The rotated condition only slightly impacts the %IoU for pedestrians and motorcyclists but more so for cyclists. Mixed conditions present a scenario with a noticeable decrease in %IoU across all classes and a resultant NDS of 35.3.
@@ IV-E Discussion of Results
Table II shows the mean average precision (mAP) of pedestrian, cyclist, and motorcyclist detection under various conditions. Under normal conditions, pedestrian detection is high at 86.3%, while cyclist detection lags at 54.7%. Motorcyclist detection is reasonably high at 77.8%, leading to an overall mAP of 72.93%. However, performance drops majorly at night and during adverse weather, with the lowest cyclist mAP at 7.5% and 8.3%, respectively. Interestingly, rotating images in normal conditions slightly reduces pedestrian and motorcyclist detection by about 1%, but notably impacts cyclist detection, dropping to 52.3%. The combination of all adverse conditions drastically reduces overall performance to a mAP of 36.7%, showcasing the challenges faced by the models in less-than-ideal scenarios.
@@
Figure 6: Figure showing mean for all three classes
@@
Figure 7: Figure with head wise attention maps for motorcyclist class.
@@
Figure 8: Figure showing mean for motorcyclist classes.
Visualizations: Figure 8 and Figure 7 shows mean and attention maps for the motorcyclist class by still capturing features from the cyclist present in the input image. The model used for testing the images has dominant weights from the motorcyclist class, thus fails to capture the features of cyclist. Similarly, Figure 6, and Figure 9 shows the results from the test which has presence of equal number of classes, but as it can be seen from the image, the model also assign weights to features represented by cyclist and pedestrians to other classes. In a summary the key takeways can be described as:
@@
  • Post-mitigation performance, which involves resampling approaches, shows improvements in both CNN (ResNet18) and ViT.
@@
Figure 9: Figure showing head-wise attention weights for all classes.
@@ V Conclusion and Future Work
Acknowledgment
The authors gratefully acknowledge funding from European Union’s Horizon 2020 Research and Innovation programme under the Marie Skłodowska Curie grant agreement No. 956090 (APROPOS), SPATIAL project under grant agreement No 101021808, and CONCORDIA project under grant agreement No 830927.